OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points (-18,-16) (-20,-19).

hero (hero):

y = mx + b -16 = -18m + b -19 = -20m + b 18m - 16 = b 20m - 19 = b b = b 18m - 16 = 20m - 19 19 - 16 = 20m - 18m 3 = 2m 3/2 = m

OpenStudy (anonymous):

I still don't understand much. What is the slope?

OpenStudy (anonymous):

so the formula is

OpenStudy (anonymous):

so the formula is

OpenStudy (anonymous):

so the formula is

OpenStudy (anonymous):

so the formula is

OpenStudy (anonymous):

so the formula is

hero (hero):

m = slope

hero (hero):

@nelly4dust, Do you understand y = mx + b?

OpenStudy (anonymous):